# Prompt-Singer: Controllable Singing-Voice-Synthesis with Natural Language Prompt

### v1

@@ 5.5 Results and Analysis
We can draw two basic conclusions from the results in table 1: 1) Generally, our models (1-10) exhibit favorable attribute controlling accuracies, with the best values being 87.7 / 86.3, 94.9 and 84.7 for the three attributes, together with competitive audio quality and melodic accuracy to non-controllable baselines (1-10 v.s. 11-13), with the best R-FFE and MOS being 0.09 and 3.90. This indicates the effectiveness of our model design on the task of controllable SVS. 2) The accuracies on volume are higher than gender and vocal range by a salient margin, with the values varying between 7.4 and 15.4 across different models. We speculate that this is because the random amplitude scaling in training allows the data with different volumes to be expanded to a large scale (somewhat similar to data augmentation), while the quantities and diversities of gender and range are limited by the training datasets. This, from one perspective, confirms that data scarcity makes learning the correlation between prompt and style attributes difficult.
@@ D.1 Objective Evaluation
Figure 4: Soft-margin accuracy curve of high vocal-range of male.
Figure 5: Soft-margin accuracy curve of medium volume.
@@
For controlling accuracies on volume and vocal range, considering that the values of generated singing may slightly deviate from the boundaries used for categorization, we adopt a soft-margin mechanism for accuracy calculation. Specifically, we take the accuracy of data falling within the correct range as 100, and calculate the accuracy with $100*\exp{(-k\epsilon)}$ for data outside the correct range, where $\epsilon$ is the error between the data value and the boundary, and $k$ is a hyper-parameter controlling the decay rate of accuracy at the margins, with larger $k$ corresponding to faster decay. We take accuracy curves of high vocal-range of male and medium volume as examples and illustrate them in Figure 4 and 5, respectively. We set $k$ to 120, 150 and 180 for high, medium and low volume, and 0.2 for vocal range accuracy.
@@ D.2 Subjective Evaluation
Our subjective evaluation tests are crowd-sourced and conducted via Amazon Mechanical Turk. For audio quality evaluation, we ask the testers to examine the audio quality and naturalness and ignore the content. For prompt-style relevance, we instruct the testers to evaluate the relevance between the natural language prompt and the singing style while ignoring the content. The testers rate scores on 1-5 Likert scales. We provide screenshots of the testing interfaces in Figure 6 and 7. Each data item is rated by 4 testers, and the testers are paid \$8 hourly.

### v2

@@ 5.5 Results and Analysis
We can draw a basic conclusion from the results in table 1: generally, our models (1-10) exhibit favorable attribute controlling accuracies, with the best values being 87.7 / 86.3, 94.9 and 84.7 for the three attributes, together with competitive audio quality and melodic accuracy to non-controllable baselines (1-10 v.s. 11-13), with the best R-FFE and MOS being 0.09 and 3.90. This indicates the effectiveness of our model design on the task of controllable SVS.
@@ D.1 Objective Evaluation
For controlling accuracies on volume and vocal range, considering that the values of generated singing may slightly deviate from the boundaries used for categorization, we adopt a soft-margin mechanism for accuracy calculation. Specifically, we take the accuracy of data falling within the correct range as 100, and calculate the accuracy with $100*\exp{(-k\epsilon)}$ for data outside the correct range, where $\epsilon$ is the error between the data value and the boundary, and $k$ is a hyper-parameter controlling the decay rate of accuracy at the margins, with larger $k$ corresponding to faster decay. We set $k$ to 10, 20 and 30 for high, medium and low volume, and 0.05 for vocal range accuracy.
@@ D.2 Subjective Evaluation
Our subjective evaluation tests are crowd-sourced and conducted via Amazon Mechanical Turk. For audio quality evaluation, we ask the testers to examine the audio quality and naturalness and ignore the content. For prompt-style relevance, we instruct the testers to evaluate the relevance between the natural language prompt and the singing style while ignoring the content. The testers rate scores on 1-5 Likert scales. We provide screenshots of the testing interfaces in Figure 4 and 5. Each data item is rated by 4 testers, and the testers are paid \$8 hourly.